# A classification result for eternal mean convex flows of finite total curvature type

### 18 Mar 2024 v1

@@ 3. Proof of theorem
Now we claim because the sets $H_{t}$ are monotone (increasing) in terms of set inclusion as $t\to-\infty$ in any ball $B(0,R)$ $H_{t}$ must fully converge to a (possibly empty, depending on $R$) set $H_{-\infty}$ in Hausdorff distance. To see this note that the space of compact subsets of a compact set (here, we consider the closure of $H_{t}$ in $B(0,R)$) is compact so that for any sequence $t_{i}\to-\infty$ $H_{t_{i}}\cap B(0,R)$ subconverges to some, possibly empty, set. Now let $t_{k},t_{\ell}\to-\infty$ be two sequences for which $H_{t_{k}}\cap B(0,R)$ Hausdorff converges to a set $S_{1}$ and $H_{t_{\ell}}\cap B(0,R)$ converges to $S_{2}$. By set monotonicity of the flow its easy to see that $S_{1}\subseteq S_{2}$ and vice versa, so they are equal implying for any sequence $t_{m}\to-\infty$ that $H_{t_{m}}$ must subconverge to a set $S$ independent of sequence. Then the full convergence claim follows by a compactness contradiction argument. As a consequence the boundaries of $H_{t}$, which are precisely $M_{t}$, must also fully converge in Hausdorff distance on bounded sets to a set which, by some abuse of notation we’ll define as $\partial H_{-\infty}$ (since a set $C$ and its closure are the same in Hausdorff distance, there is some ambiguity when speaking of the boundary of $\partial H_{\infty}$). By the assumed curvature bound, namely that in balls of fixed radius $M_{t}$ can be written as a union of graphs of uniformly bounded gradient, we see the support of $M_{-\infty,t}$ must be precisely $\partial H_{-\infty}$.
Now, because the $M_{t}$ are all embedded we see that $\partial H_{-\infty}$, in some suitably small ball $B$, can be locally written as the union of smooth graphs which at worst intersect tangentially. The flow $M_{-\infty,t}$ from the paragraph above is also represented locally by the union of these graphs, although one might imagine with different multiplicities. At any rate because the flow $M_{-\infty,t}$, which is smoothly varying, is precisely $\partial H_{-\infty}$ for all times these graphs must each be minimal. Then by the maximum principle these graphs must either agree or be disjoint so $\partial H_{-\infty}$ must actually be an embedded minimal surface.
With this in mind one can see for fixed $t$ that we have full/true convergence of $M_{t+s}$ to the backwards limit. In this case we already know that $\partial H_{-\infty}$ is smooth and embedded we see in combination with the uniform curvature bound we may apply the first part of the argument of the lemma below, lemma 3.6, to see the convergence of $M_{t+s}$ to $\partial H_{-\infty}$ must be with multiplicity one – the gist is that one may use the set monotonicity of the flow to see if the convergence was greater than 2 somewhere then $M_{t}$ could not intersect $\partial H_{-\infty}$ and there would have to be a sheet on either side of $\partial H_{-\infty}$, so $\partial H_{-\infty}$ would disconnect $M_{t}$ giving a contradiction (the flow is assumed to be connected). In particular, for any sequence $t_{i}\to-\infty$ for which $M_{t_{i}}$ converges in the sense of varifolds the limit varifold must be the same ($\partial H_{-\infty}$ with multiplicity one). With this in mind its reasonable to denote then $\partial H_{-\infty}$ $M_{-\infty}$, and a compactness contradiction argument (using Allard’s compactness theorem via the entropy bound and finite total curvature type assumption) gives that the convergence $M_{t}\to M_{-\infty}$ is full in the sense of varifolds.
Now, let’s suppose that the curvature of $M_{t}$ is not uniformly pointwise bounded but just satisfies the conditions of the theorem. By the finite total curvature type condition and entropy bound we may employ Brakke’s compactness theorem with the flows $M_{t+s_{i}}$ where $s_{i}\to-\infty$ to obtain a subsequential limit Brakke flow which we’ll denote here by $M_{-\infty,t}$, apriori dependent on the choice of sequence taken. Of course (as a weaker statement) for $t$ fixed under our assumptions by Allard’s compactness theorem for any sequence $s_{i}\to-\infty$ we may extract a subsequence $s_{k}\to-\infty$ such that $M_{t+s_{k}}$ varifold converges, and it follows from the reasoning in the paragraph above that for any sequence $s_{k}\to-\infty$ for which $M_{t+s_{k}}$ converge as varifolds that the support of the limit varifold $\nu$ is contained in $\partial H_{-\infty}$, which by abuse of terminology as above we define as the Hausdorff limit of $M_{t}$. Similar to above we want to study next to what extent is the limit $\nu$ dependent on the sequence $s_{k}$.
To proceed it’s helpful to have a more concrete handle on the nature of the convergence $M_{t}$ to a limit varifold. Since we assume $M_{t}$ has finite entropy and satisfies part (b) of the finite total curvature type condition we draw inspiration from Ilmanen’s analysis in [16] to do so – see also [24] for a nice summary. Sketching this out, considering the measures $\sigma_{t}=|A|^{2}\mathcal{H}^{2}\llcorner M_{t}$ by the finite total curvature type condition for a sequence $t+s_{k}\to-\infty$ we may take a subsequence $t+s_{\ell}\to-\infty$ for which these radon measures $\sigma_{t+s_{\ell}}$ converge to a limit measure $\sigma$. By Simon’s sheeting theorem in balls $B$ where $\sigma_{t}$ is sufficiently small (i.e. where the intergral curvature doesn’t concentrate) we can locally write $M_{t}\cap B$ as a union of embedded discs with controlled area. With this in mind, suppose that $t+s_{k}\to-\infty$ is a sequence for which as varifolds $M_{t}$ converge to some measure $\nu$. Denoting the support of $\nu$ by $N$, then by the finite total curvature type assumption and the sheeting theorem it follows there is a subsequence $t+s_{\ell}$ as above so that in each ball $B(0,R)$ there will be a finite (possibly empty) set of points $Q=Q(R)$ for which the convergence $M_{t}\to N$ in $B(0,R)$ will be multisheeted over $N\setminus Q$. In particular, writing the set $Q=\{p_{1},\ldots,p_{k}\}$ we have there are $r_{s_{\ell}}\to 0$ for which in the complement of $\cup_{i=1}^{k}B(p_{i},r_{i}(s_{\ell}))$ $M_{t+s_{\ell}}\cap B(0,R)$ is the union of disjoint surfaces “sheeted” over $N$, which may be connected/bridged within the balls.
With the paragraph above in hand and that $\partial H_{-\infty}$ is the Hausdorff limit of $M_{t}$ we next show that the/a limit measure $\nu$ as described above must in fact equal $\partial H_{-\infty}$ up measure zero inspired by the proof of corollary 5.3 of [24]. Supposing this is not the case, then of course if we further pass to the subsequence $M_{t+s_{\ell}}$ defined above the limit of sequence will also not have support equal to $\partial H_{-\infty}$ (indeed, the limit measure is still $\nu$), so it suffices to consider this sequence to gain a contradiction. Working towards a contradiction, suppose then $p\in\partial H_{-\infty}$ is not in the support $N$ of $\nu$. Then there is a $B(p,r)$ for which $N$ has measure zero in $B(p,r)$. Because $\partial H_{-\infty}$ is the Hausdorff limit of $M_{t}$ and the convergence is full for $t$ sufficiently large $M_{t}\cap B(p,r/2)$ is nonempty. Considering all this within a large fixed ball $B(0,R)$, from the paragraph above there is a finite set of points $Q$ away from which the convergence of $M_{t+s_{\ell}}$ is multisheeted, so in particular for $t+s_{\ell}$ sufficiently negative there is a finite set of points $\{p_{1},\ldots,p_{k}\}$ so that $M_{t+s_{\ell}}\cap(\cup_{i}B(p_{i},r/1000))^{c}$ is given by a union of disjoint embedded surfaces. Because for $t+s_{\ell}$ sufficiently negative $M_{t+s_{\ell}}$ passes through $B(p,r/2)$ we then get a contradiction, because as a consequence of this and of the sheeting theorem (the discs in the sheeting theorem have a lower area bound as well as an upper bound) the measure of $M_{t+s_{\ell}}\cap B(p,r)$ is uniformly bounded below for very negative $t+s_{\ell}$ giving a contradiction.
Now, a technical issue is that the multiplicity of convergence here may depend on the sequence $s_{k}$. Above we showed first that $\partial H_{-\infty}$ was minimal and then showed the multiplicity of convergence was independent of sequence, but how we will proceed here the steps are reversed. Note since $M_{t}$ is embedded the number of sheets in the sheeting theorem must agree in overlaps of balls where it applies, so since we can apply it away from a discrete set of points the multiplicity of convergence must be constant on connected components of $\partial H_{-\infty}$ for a given sequence $s_{k}$. By the uniform entropy bound, this number is uniformly bounded above independent of the sequence. Suppose on a connected component of $\partial H_{-\infty}$ the convergence is with multiplicity $a$ along $s_{k}$ and $b$ along say the sequence $s_{k}^{\prime}$ with $a\neq b$. Intertwining the sequences to get a sequence $s_{\ell}$, we see that by the smoothness of the flow there must be times $t_{k}$ between $t+s_{\ell}$ and $t+s_{\ell+1}$ for which the flow is not sheeted away from a set of small balls over $\partial H_{-\infty}\cap B(0,R)$. Taking a subsequence of this as above (using the sheeting theorem, etc.) we gain a contradiction, so the multiplicity of convergence must be independent of converging sequenece. Then by another compactness contradiction argument with Allard’s compactness theorem we can see the varifold convergence of $M_{t+s}$ is full/true in $s\to-\infty$. Considering the defintion of Brakke flow convergence this gives the Brakke flow $M_{-\infty,t}$ defined above is independent of the sequence $s_{i}\to-\infty$ choosen, implying $M_{-\infty,t}$ $M_{-\infty,t+\delta}$ are the same for any $\delta$. Hence we find again that $M_{-\infty,t}$ is time independent. As discussed in chapter 5 of [25], this implies that $M_{t}$ is stationary or so that $|H|=0$ a.e. on $M_{-\infty,t}$. We write this stationary set as $M_{-\infty}$ like before.
Because $|H|=0$ a.e. on $M_{-\infty}$, so that its minimal, when considering the measure convergence on just one of these sheets in a ball where $\sigma$ doesn’t concentrate then by applying Allard’s regularity theorem we get that $M_{-\infty}$ has $C^{1,\alpha}$, and so smooth by regularity theory for minimal surfaces in balls, support. This reasoning applies in balls with centers in all of $M_{-\infty}$ except potentially at a discrete set of points where the measures concentrate – these correspond to places where the sheets are potentially bridged. These can be pieced together to realize $M_{-\infty}$ globally as an immersed minimal surface away from the points mentioned. Gulliver’s removable singularity theorem then give $M_{-\infty}$ is a smooth minimal surface which can be seen to be embedded because the flow $M_{t}$ is using the maximum principle. As a result its also complete. $\square$
Before using this to say something more about $M_{t}$ first we refine the topology this convergence is in, where actually the information about the topology of $M_{t}$ we’ve already gained comes into play in a minor (possibly inessential) way. Below we strongly use the assumption that $M_{t}$ is connected: note for embedded minimal surfaces $\Sigma$ of finite total curvature one create ancient flows which propogate into both bounded domains of $\Sigma$, and considering these as one single (disconnected) ancient mean curvature flow they will flow back to $\Sigma$ as $t\to-\infty$ with multiplicity two.
@@
The convergence of $M_{t}\to M_{-\infty}$ is smooth with multiplicity one, and thus the components of $M_{-\infty}$ are smooth complete minimal surfaces of finite total curvature.
@@
We suppose that $M_{-\infty}$ is nonempty, or else there is nothing to do. We note by Brakke’s regularity theorem and Shi’s estimates the convergence of $M_{t}\to M_{-\infty}$ will be smooth if we can show the multiplicity of the convergence in the argument above is one by a compactness contradiction argument – an easy way to see we have the necessary density bounds to apply the regularity theorem is by the one sided minimizing property for mean convex flows (again, [27]) used with small tubular neighborhoods of $M_{-\infty}\cap B(0,R)$. By the finite total curvature type assumption this will also imply $M_{-\infty}$ has bounded total curvature, and as above the limit will be complete.
@@
As pointed out in the proof above as a consequence of Simon’s sheeting theorem we have that for any sequence $s_{i}\to-\infty$, after potentially taking a subsequence, most, if not all of $M_{s_{i}}$ can be written as the union of some number of disjoint sheets over $M_{-\infty}$ in a given ball $B(0,R)$ for $s_{i}$ sufficiently negative – by “most” here we mean away from a set (for $t$ sufficiently negative) of small disjoint balls within which there may be “bridges” connecting the sheets. The number of sheets must be finite by the finite entropy assumption, of course, and the lemma above gives that the multiplicity of convergence is independent of such a sequence. Our main goal is to gain a contradiction if there is such a sequence $s_{i}\to-\infty$ for which $M_{s_{i}}\to M_{-\infty}$ with multiplicity $m$ two or greater.
@@
We first claim there indeed must be bridges connecting at least some of these sheets where $R$ is large enough, corresponding to where the curvature concentrates in the limit – it isn’t immediate there should be any despite $M_{t}$ being connected because we are in the noncompact case and so apriori where the sheets “connect/wrap around” could tend to spatial infinity in the backwards limit, and also where the curvature concentrates there could just be “pimples” instead of actual bridges. More precisely what we are claiming is that if $R$ is sufficiently large there are a finite (nonempty) set of balls $B(p_{1}(s_{i}),r_{1}(s_{i})),\ldots B(p_{N(s_{i})}(s_{i}),r_{N(s_{i})}(s_{i}))$ in $B(0,R)$ for which ($M_{s_{i}}\cap B(0,R))\setminus\{\cup_{j=1}^{N(s_{i})}B(p_{j}(s_{i}),r_{j}(s_{i}))\}$ is a union of $m(s_{i})$ disjoint smooth manifolds with boundary, that $M_{s_{i}}\cap B(0,R)$ has $1\leq c<m(s_{i})$ connected components, and the maximum radius of the balls above tends to zero as $s_{i}\to-\infty$. Here by $N(s_{i})$ and $m(s_{i})$ we denote the number of balls $N(s_{i})$ and number of sheets $m(s_{i})$ respectively. These numbers must eventually stabilize by the choice of sequence, but imaginably it could be the case that $m(s_{i})\neq m$ for $s_{i}$ relatively positive.

### 14 Mar 2026 v2

@@ 3. Proof of theorem
First note that the space of compact subsets of a compact set (here, we consider the closure of $H_{t}$ in $B(0,R)$) is compact so that for any sequence $t_{i}\to$-$\infty$ $H_{t_{i}}\cap B(0,R)$ subconverges to some, possibly empty, set. Now let $t_{k},t_{\ell}\to$-$\infty$ be two sequences for which $H_{t_{k}}\cap B(0,R)$ Hausdorff converges to a set $S_{1}$ and $H_{t_{\ell}}\cap B(0,R)$ converges to $S_{2}$. By set monotonicity of the flow it’s easy to see that $S_{1}\subseteq S_{2}$ and vice versa, so they are equal implying for any sequence $t_{m}\to$-$\infty$ that $H_{t_{m}}$ must subconverge to a set $S$ independent of sequence. Then the full convergence claim follows by a compactness contradiction argument. As a consequence the boundaries of $H_{t}$, which are precisely $M_{t}$, must also fully converge in Hausdorff distance on bounded sets to a set which, by some abuse of notation we’ll define as $\partial H_{-\infty}$ (since a set $C$ and its closure are the same in Hausdorff distance, there is potentially some ambiguity when speaking of the boundary of $\partial H_{\infty}$). $\square$
We next refine the geometry of this limit; For the sake of exposition first let’s show a result in the case $M_{t}$ has pointwise uniformly bounded curvature, by which we mean $|A|^{2}(p,t)<C$ for some fixed constant $C$ independent of $p$ and tt– this is indeed satisfied by the eternal flow out of a catenoid mentioned in the introduction, but a priori this isn’t immediate. Precisely, let’s first show:
@@
Supposing in addition to the assumptions of Theorem 1.1 above that $|A|^{2}(p,t)$ is bounded by a uniform constant $C$, the set $\lim\limits_{t\to-\infty}M_{t}$ is either empty or diffeomorphic to a disjoint union of smooth embedded complete minimal surfaces, where the convergence is full in the sense of varifolds for any given fixed bounded domain.
@@
Note that by the pointwise curvature bound and Shi’s estimates that $|\nabla^{k}A|$ for any $k\geq 1$ is bounded uniformly along $M_{t}$ for all times $t<0$. This gives by the entropy bound and Arzela–Ascoli subsequential convergence of the flows $M_{t+s_{i}}$ for a sequence $s_{i}\to$-$\infty$ in $C^{k}$ norm (a priori, possibly multisheeted) for any $k\geq 0$ on bounded sets. The limit flow will be smooth (but potentially only immersed) and evolve by the classical mean curvature flow. Naturally denoting this limit flow by $M_{-\infty,t}$, we see that a priori it may depend on the sequence of $s_{i}$ but actually because the Hausdorff convergence in lemma 3.5 above is full we see this is not the case and in fact by the assumed curvature bound, namely that in balls of fixed radius $M_{t}$ can be written as a union of graphs of uniformly bounded gradient, we see the support of $M_{-\infty,t}$ must be precisely $\partial H_{-\infty}$.
Now, because the $M_{t}$ are all embedded we see that $\partial H_{-\infty}$, in some suitably small ball $B$, can be locally written as the union of smooth graphs which at worst intersect tangentially. The flow $M_{-\infty,t}$ is also represented locally by the union of these graphs, although one might imagine with different multiplicities. At any rate because the flow $M_{-\infty,t}$, which is smoothly varying, is precisely $\partial H_{-\infty}$ for all times these graphs must each be minimal. Then by the maximum principle these graphs must either agree or be disjoint so $\partial H_{-\infty}$ must actually be an embedded minimal surface.
We next claim that for fixed $t$ as $s\to$-$\infty$ that the convergence of $M_{t+s}$ to $\partial H_{-\infty}$ must be with multiplicity one – by the curvature bound assumption we see the convergence must be locally graphical but with multiple connected components. To see this, by the set monotonicity of the flow if the backwards convergence was greater than 2 at some point of $\partial H_{-\infty}$ then $M_{t}$ could not intersect $\partial H_{-\infty}$ and there would have to be a sheet on either side of $\partial H_{-\infty}$, so $\partial H_{-\infty}$ would disconnect $M_{t}$ giving a contradiction (the flow is assumed to be connected). In particular, for any sequence $t_{i}\to$-$\infty$ for which $M_{t_{i}}$ converges in the sense of varifolds the limit varifold must be the same ($\partial H_{-\infty}$ with multiplicity one). With this in mind it’s reasonable to denote then $\partial H_{-\infty}=M_{-\infty}$, and a compactness–contradiction argument (using Allard’s compactness theorem via the entropy bound and finite total curvature type assumption) gives that the convergence $M_{t}\to M_{-\infty}$ is full in the sense of varifolds. $\square$
In fact, because the backwards convergence above is with multiplicity one it is not too hard to see it is smooth and we discuss this more below. To deal with the more general case we employ the finite total curvature type assumption along with Simon’s sheeting theorem as employed by Ilmanen in [16].
Lemma 3.7.
The set $\lim\limits_{t\to-\infty}M_{t}$ is either empty or diffeomorphic to a disjoint union of smooth embedded complete minimal surfaces, where the convergence is locally in the sense of varifolds possibly with multiplicity.
Proof:
By the finite total curvature type condition and entropy bound we may employ Brakke’s compactness theorem with the flows $M_{t+s_{i}}$ where $s_{i}\to$-$\infty$ to obtain a subsequential limit Brakke flow which we’ll denote here by $M_{-\infty,t}$, a priori dependent on the choice of sequence taken. Of course (as a weaker statement) for $t$ fixed under our assumptions by Allard’s compactness theorem for any sequence $s_{i}\to$-$\infty$ we may extract a subsequence $s_{k}\to$-$\infty$ such that $M_{t+s_{k}}$ varifold converges, and it follows from lemma 3.5 for any sequence $s_{k}\to$-$\infty$ for which $M_{t+s_{k}}$ converge as varifolds that the support of the limit varifold $\nu$ is contained in $\partial H_{-\infty}$, which by abuse of terminology as above we define as the Hausdorff limit of $M_{t}$. Similar to above we want to study next to what extent is the limit $\nu$ dependent on the sequence $s_{k}$.
To proceed it’s helpful to have a more concrete handle on the nature of the convergence $M_{t}$ to a limit varifold. Since we assume $M_{t}$ has finite entropy and satisfies part (b) of the finite total curvature type condition we draw inspiration from Ilmanen’s analysis in [16] to do so – see also [25] for a nice summary. Sketching this out, considering the measures $\sigma_{t}=|A|^{2}\mathcal{H}^{2}\llcorner M_{t}$, where $\mathcal{H}^{2}$ is the 2–dimensional Hausdorff measure, by the finite total curvature type condition for a sequence $t+s_{k}\to$-$\infty$ we may take a subsequence $t+s_{\ell}\to$-$\infty$ for which these radon measures $\sigma_{t+s_{\ell}}$ converge to a limit measure $\sigma$. By Simon’s sheeting theorem in balls $B$ where $\sigma_{t}$ is sufficiently small (i.e. where the integral curvature doesn’t concentrate) we can locally write $M_{t}\cap B$ as a union of embedded discs with controlled area. With this in mind, suppose that $t+s_{k}\to$-$\infty$ is a sequence for which as varifolds $M_{t}$ converge to some measure $\nu$. Denoting the support of $\nu$ by $N$, then by the finite total curvature type assumption and the sheeting theorem it follows there is a subsequence $t+s_{\ell}$ as above so that in each ball $B(0,R)$ there will be a finite (possibly empty) set of points $Q=Q(R)$ for which the convergence $M_{t}\to N$ in $B(0,R)$ will be multisheeted over $N\setminus Q$. In particular, writing the set $Q=\{p_{1},\ldots,p_{k}\}$ we have there are $r_{s_{\ell}}\to 0$ for which in the complement of $\cup_{i=1}^{k}B(p_{i},r_{i}(s_{\ell}))$ $M_{t+s_{\ell}}\cap B(0,R)$ is the union of disjoint surfaces “sheeted” over $N$, which may be connected/bridged within the balls.
We next show that the support $N$ of the/a limit measure $\nu$ as described above must in fact equal $\partial H_{-\infty}$ up measure zero inspired by the proof of corollary 5.3 of [25]. Supposing this is not the case, then of course if we further pass to the subsequence $M_{t+s_{\ell}}$ defined above the limit of sequence will also not have support equal to $\partial H_{-\infty}$ (indeed, the limit measure is still $\nu$), so it suffices to consider this sequence to gain a contradiction. Working towards a contradiction, suppose then $p\in\partial H_{-\infty}$ is not in the support $N$ of $\nu$. Then there is a $B(p,r)$ for which $N$ has measure zero in $B(p,r)$. Because $\partial H_{-\infty}$ is the Hausdorff limit of $M_{t}$ and the convergence is full for $t$ sufficiently large $M_{t}\cap B(p,r/2)$ is nonempty. Considering all this within a large fixed ball $B(0,R)$, from the paragraph above there is a finite set of points $Q$ away from which the convergence of $M_{t+s_{\ell}}$ is multisheeted, so in particular for $t+s_{\ell}$ sufficiently negative there is a finite set of points $\{p_{1},\ldots,p_{k}\}$ so that $M_{t+s_{\ell}}\cap(\cup_{i}B(p_{i},r/1000))^{c}$ is given by a union of disjoint embedded surfaces. Because for $t+s_{\ell}$ sufficiently negative $M_{t+s_{\ell}}$ passes through $B(p,r/2)$ we then get a contradiction, because as a consequence of this and of the sheeting theorem (the discs in the sheeting theorem have a lower area bound as well as an upper bound) the measure of $M_{t+s_{\ell}}\cap B(p,r)$ is uniformly bounded below for very negative $t+s_{\ell}$ giving a contradiction.
Now, a technical issue is that the multiplicity of convergence here may depend on the sequence $s_{k}$. In the expository lemma above, lemma 3.6, we first showed that $\partial H_{-\infty}$ was minimal and then showed the multiplicity of convergence was independent of sequence, but how we will proceed here the steps are reversed and this point on multiplicity we’ll soon see is important. Note since $M_{t}$ is embedded the number of sheets in the sheeting theorem must agree in overlaps of balls where it applies, so since we can apply it away from a discrete set of points the multiplicity of convergence must be constant on connected components of $\partial H_{-\infty}$ for a given sequence $s_{k}$. By the uniform entropy bound, this number is uniformly bounded above independent of the sequence. Suppose on a connected component of $\partial H_{-\infty}$ the convergence is with multiplicity $a$ along $s_{k}$ and $b$ along say the sequence $s_{k}^{\prime}$ with $a\neq b$. Intertwining the sequences to get a sequence $s_{\ell}$, we see that by the smoothness of the flow there must be times $t_{k}$ between $t+s_{\ell}$ and $t+s_{\ell+1}$ for which the flow is not sheeted away from a set of small balls over $\partial H_{-\infty}\cap B(0,R)$. Taking a subsequence of this as above (using the sheeting theorem, etc.) we gain a contradiction, so the multiplicity of convergence must be independent of converging sequence.
As a consequence, by another compactness contradiction argument with Allard’s compactness theorem, we can see the varifold convergence of $M_{t+s}$ is full/true in $s\to$-$\infty$ (again, within any given fixed bounded domain of ${\mathbb{R}}^{3}$). Considering the definition of Brakke flow convergence this gives the Brakke flow $M_{-\infty,t}$ defined above is independent of the sequence $s_{i}\to$-$\infty$ chosen, implying $M_{-\infty,t}$ $M_{-\infty,t+\delta}$ are the same for any $\delta$. Hence we find again that $M_{-\infty,t}$ is time independent. As discussed in chapter 5 of [26], this implies that $M_{t}$ is stationary or so that $|H|=0$ a.e. on $M_{-\infty,t}$. We write this stationary set as $M_{-\infty}$ like in the expository lemma.
Because $|H|=0$ a.e. on $M_{-\infty}$, so that it’s minimal, when considering the measure convergence on just one of these sheets in a ball where $\sigma$ doesn’t concentrate then by applying Allard’s regularity theorem we get that $M_{-\infty}$ has $C^{1,\alpha}$, and so smooth by regularity theory for minimal surfaces in balls, support. This reasoning applies in balls with centers in all of $M_{-\infty}$ except potentially at a discrete set of points where the measures concentrate – these correspond to places where the sheets are potentially bridged. These can be pieced together to realize $M_{-\infty}$ globally as an immersed minimal surface away from the points mentioned. Gulliver’s removable singularity theorem then give $M_{-\infty}$ is a smooth minimal surface which can be seen to be embedded because the flow $M_{t}$ is using the maximum principle. As a result it is also complete. $\square$
Before using this to say something more about $M_{t}$ (for finite times) first we refine the topology this convergence is in, where actually the information about the topology of $M_{t}$ we’ve already gained comes into play in a minor (possibly inessential) way. Below we strongly use the assumption that $M_{t}$ is connected: note for embedded minimal surfaces $\Sigma$ of finite total curvature one create ancient flows which propagate into both bounded domains of $\Sigma$, and considering these as one single (disconnected) ancient mean curvature flow they will flow back to $\Sigma$ as $t\to$-$\infty$ with multiplicity two.
Lemma 3.8.
The convergence of $M_{t}\to M_{-\infty}$ is locally smooth with multiplicity one, and thus the components of $M_{-\infty}$ are smooth complete minimal surfaces of finite total curvature.
Proof:
@@
We suppose that $M_{-\infty}$ is nonempty, or else there is nothing to do. We note by Brakke’s regularity theorem and Shi’s estimates the convergence of $M_{t}\to M_{-\infty}$ will be in $C^{\infty}_{loc}$ if we can show the multiplicity of the convergence in the argument above is one by a compactness contradiction argument – an easy way to see we have the necessary density bounds to apply the regularity theorem is by the one sided minimizing property for mean convex flows (again, [28]) used with small tubular neighborhoods of $M_{-\infty}\cap B(0,R)$. By the finite total curvature type assumption this will also imply $M_{-\infty}$ has bounded total curvature, and as above the limit will be complete.
@@
As pointed out in the proof above as a consequence of Simon’s sheeting theorem we have that for any sequence $s_{i}\to$-$\infty$, after potentially taking a subsequence, most, if not all of $M_{s_{i}}$ can be written as the union of some number of disjoint sheets over $M_{-\infty}$ in a given ball $B(0,R)$ for $s_{i}$ sufficiently negative – by “most” here we mean away from a set (for $t$ sufficiently negative) of small disjoint balls within which there may be “bridges” connecting the sheets. The number of sheets must be finite by the finite entropy assumption, of course, and the lemma above gives that the multiplicity of convergence is independent of such a sequence. Our main goal is to gain a contradiction if there is such a sequence $s_{i}\to$-$\infty$ for which $M_{s_{i}}\to M_{-\infty}$ with multiplicity $m$ two or greater. Because dealing with each of the possible cases is relatively lengthy for the reader’s convenience below we break the argument into a number of highlighted subclaims:
@@
Subclaim 1: there must be "bridges" joining sheets. We first claim there indeed must be bridges connecting at least some of these sheets where $R$ is large enough, corresponding to where the curvature concentrates in the limit – it isn’t immediate there should be any despite $M_{t}$ being connected because we are in the noncompact case and so a priori where the sheets “connect/wrap around” could tend to spatial infinity in the backwards limit, and also where the curvature concentrates there could just be “pimples” instead of actual bridges. More precisely what we are claiming is that if $R$ is sufficiently large there are a finite (nonempty) set of balls $B(p_{1}(s_{i}),r_{1}(s_{i})),\ldots B(p_{N(s_{i})}(s_{i}),r_{N(s_{i})}(s_{i}))$ in $B(0,R)$ for which ($M_{s_{i}}\cap B(0,R))\setminus\{\cup_{j=1}^{N(s_{i})}B(p_{j}(s_{i}),r_{j}(s_{i}))\}$ is a union of $m(s_{i})$ disjoint smooth manifolds with boundary, that $M_{s_{i}}\cap B(0,R)$ has $1\leq c<m(s_{i})$ connected components, and the maximum radius of the balls above tends to zero as $s_{i}\to$-$\infty$. Here by $N(s_{i})$ and $m(s_{i})$ we denote the number of balls $N(s_{i})$ and number of sheets $m(s_{i})$ respectively. These numbers must eventually stabilize by the choice of sequence, but imaginably it could be the case that $m(s_{i})\neq m$ for $s_{i}$ relatively positive.